Determine whether each of the equations is exact. If it is exact, find the solution.

  1. (dy/dx)=-(ax-by/bx-cy)
  2. (2x+4y)+(2x-2y)y'=0
  3. (3x2-2xy+2)+(6y2-x2+3)y'=0
  4. (ycosx+2xey)+(sinx+x2ey-1)y'=0
1
Expert's answer
2021-11-28T17:15:30-0500

"\text{Differential equation,}\\
M(x,y)dx+N(x,y)=0\\
\text{then,
if partial derivative of M w.r.t y is equal to the partial derivative of N w.r.t x. ---(1)}\\
1.\\
\frac{dy}{dx}=-\frac{ax-by}{bx-cy}\\
(ax-by)dx+(bx-cy)dy=0\\
Now,\\
M_y=-b\\
N_x=b\\
\text{This is not exact.}\\
2.\\
(2x+4y)dx+(2x-2y)dy=0\\
Now,\\
M_y=4\\
N_x=2\\
\text{This is not exact.}\\
3.\\
(3x^2-2xy+2)dx+(6y^2-x^2+3)dy=0\\
Now,\\
M_y=-2x\\
N_x=-2x\\
\text{This is an exact differential equation.}\\
\text{Then the solution is given by,}\\
\int(3x^2-2xy+2)dx+\int(6y^2-x^2+3)dy=c\\
x^3-x^2y+2x+2y^3+3y=c\\
4.\\
(ycosx+2xe^y)dx+(sinx+x^2e^y-1)dy=0
Now,\\
M_y=cosx+2xe^y\\
N_x=cosx+2xe^y\\
\text{This is an exact differential equation.}\\
\text{Then the solution is given by,}\\
\int(ycosx+2xe^y)dx+\int(sinx+x^2e^y-1)dy=c\\
ysinx+x^2e^y-y=c"